# MCPDepth: Omnidirectional Depth Estimation via Stereo Matching from Multi-Cylindrical Panoramas

### v1

@@ 3.2 Stereo Matching
Figure 4: (a) displays the circular attention module in the stereo matching network. The number of attention heads is 8. (b) represents our attention applied along the vertical axis. $\bigoplus$ denotes element-wise sum. $\bigotimes$ denotes matrix multiplication. Blue boxes are $1\times 1$ convolution and orange boxes are relative positional encoding.
@@
Circular Attention To overcome vertical axis distortion and capture the circular $360^{\circ}$ features, we introduce a circular attention module. Conventional CNNs have limited receptive fields, which is restrictive for $360^{\circ}$ FoV panoramas. The circular attention module, placed between feature extraction and cost volume construction is flexible and can be easily integrated since it maintains the input dimension. Besides, it only calculates the relations along the vertical axis, conserving more training resources compared to global self-attention approaches. Fig. 4 (a) demonstrates our circular attention module.
@@
where $\mathcal{N}_{1\times m}(o)$ is the local $1\times m$ region centered around location $o=(i,j)$. $r^{q}_{p-o}\in\mathbb{R}^{d_{q}}$ is the learnable relative positional encoding for queries and the inner product $q^{T}_{o}r^{q}_{p-o}$ measures the compatibility from location $p$ to location $o$. Similarly, the learnable vectors $r^{k}_{p-o}\in\mathbb{R}^{d_{q}}$ and $r^{v}_{p-o}\in\mathbb{R}^{d_{out}}$ are positional encodings for keys and values. Our circular attention module reduces the computation to $(\mathcal{O}(hwm))$.
@@ 4.1 Datasets
We train and evaluate our framework on Deep360 [20] and 3D60 [61], which include outdoor synthetic and indoor real scenes. We evaluate both stereo matching and depth estimation tasks on them. For Deep360, 4 $360^{\circ}$ cameras are arranged on a horizontal plane to form a square. Panoramas from all 4 views are used to validate the performance of our framework. We use 6 pairs of stereo images for training and testing. For 3D60, 3 $360^{\circ}$ cameras are arranged on a vertical plane to form an equilateral right triangle. Panoramas from two of three views are used to validate the performance of our framework. The resolutions of Deep360 and 3D60 are $H\times W=1024\times 512$ and $H\times W=512\times 256$, respectively.
@@ 4.2 Evaluation Metrics
For a fair comparison, we follow MODE [20] and evaluate stereo matching performance using MAE (mean absolute error), RMSE (root mean square error), Px1,3,5 (percentage of outliers with pixel error $>$ 1, 3, 5), D1 [30] (percentage of outliers with pixel error $>$ 3 and $>$ 5%). We evaluate depth estimation performance using MAE, RMSE, AbsRel (absolute relative error), SqRel (square relative error), SILog [8](scale-invariant logarithmic error), $\delta 1,2,3$ [17] (accuracy with threshold that $\max(\tfrac{\hat{y}}{y^{\star}},\tfrac{y^{\star}}{\hat{y}})<1.25,1.25^{2},1.25^{3}$).
@@ 4.3 Implementation Details
We use the nearest-neighbor interpolation along the horizontal axis to generate cylindrical disparity ground truth from spherical angular disparity and bi-linear interpolation to generate cylindrical panoramas from spherical ones.
In the stereo matching stage, cylindrical panoramas have a $360^{\circ}$ vertical FoV and a horizontal FoV of less than $180^{\circ}$. We evaluate the central part of disparity maps in the Cassini domain with horizontal FoV = $2*\arctan(\pi/2)\approx 105^{\circ}$ for both datasets as shown in Fig. 1 (d). This horizontal FoV can keep the size of cylindrical panoramas equal to spherical panoramas. Note that only when the horizontal FoV is greater than $\pi/4$ can the invisible areas of a certain viewpoint be compensated by other viewpoints. In the fusion stage, we evaluate the entire omnidirectional depth map with $360^{\circ}$ horizontal FoV and $180^{\circ}$ vertical FoV.
4.4 Results
Training on Perspective Images and Testing on Panoramas The pre-trained models of PSMNet [3] and IGEV-Stereo [54] are trained on Sceneflow [27], which contain only perspective images. CREStereo [18], trained on mixed datasets, exhibits better generalization. The performance of stereo matching with different projections on Deep360 is shown in Tab. 1. Acquiring panoramas and their depth ground truth is difficult, the experimental results provide the potential to apply stereo-matching models trained on perspective images to cylindrical panoramas.
@@ 4.4 Results
Table 2: Quantitative results of stereo matching methods on Deep360 and 3D60 test datasets.
Dataset Method Projection MAE RMSE Px1 (%) Px3 (%) Px5 (%) D1 (%)
Deep360 AANet [55] Cassini 0.3427 1.5703 5.2050 2.1515 1.2847 1.9817
360SD-Net [46] Cassini 0.5262 1.6459 3.8794 1.3389 0.8425 1.2989
PSMNet [3] Cassini 0.2703 1.4790 3.3556 1.1979 0.7538 1.1708
MODE [20] Cassini 0.2309 1.4014 2.8801 1.0488 0.6562 1.0326
Ours Cylindrical 0.2112 1.3903 2.5713 1.0009 0.6376 0.9828
3D60 MODE [20] Cassini 0.2258 0.5265 2.9441 0.6482 0.2978 0.6478
Ours Cylindrical 0.1773 0.4654 2.2298 0.5282 0.2564 0.5279
For omnidirectional depth estimation, we compare our method with other multi-view omnidirectional depth estimation methods including UNiFuse [15], CSDNet [19], 360SD-Net [46], OmniMVS [50], and MODE [20]. We report the results from MODE. Tab. 3 shows that our method achieves an 18.8% MAE reduction on Deep360 and 19.9% on 3D60 from the best-published result.
@@
Dataset Method MAE($\downarrow$) RMSE($\downarrow$) AbsRel($\downarrow$) SqRel($\downarrow$) SILog($\downarrow$) $\delta 1(\%\uparrow)$ $\delta 2(\%\uparrow)$ $\delta 3(\%\uparrow)$
Deep360 UniFuse [15] 3.9193 28.8475 0.0546 0.3125 0.1508 96.0269 98.2679 98.9909
CSDNet [19] 6.6548 36.5526 0.1553 1.7898 0.2475 86.0836 95.1589 97.7562
360SD-Net [46] 11.2643 66.5789 0.0609 0.5973 0.2438 94.8594 97.2050 98.1038
OmniMVS [50] 8.8865 59.3043 0.1073 2.9071 0.2434 94.9611 97.5495 98.2851
MODE [20] 3.2483 24.9391 0.0365 0.0789 0.1104 97.9636 99.0987 99.4683
Ours 2.6384 21.6692 0.0304 0.1153 0.1033 98.2557 99.2101 99.5227
3D60 UniFuse [15] 0.1868 0.3947 0.0799 0.0246 0.1126 93.2860 98.4839 99.4828
CSDNet [19] 0.2067 0.4225 0.0908 0.0241 0.1273 91.9537 98.3936 99.5109
360SD-Net [46] 0.0762 0.2639 0.0300 0.0117 1.4578 97.6751 98.6603 99.0417
MODE [20] 0.0713 0.2631 0.0224 0.0031 0.0512 99.1283 99.7847 99.9250
Ours 0.0571 0.1903 0.0199 0.0027 0.0401 99.3933 99.8506 99.9418
Fig. 5 illustrates our method’s superior depth estimation on Deep360, with better handling of severe distortions, finer object details, and edges between foreground and background compared to MODE. More visual comparisons on Deep360 and 3D60 datasets are available in the supplementary material.
@@
Performance on Real Scenarios We evaluate our models on real-world fisheye image pairs with a FoV of $189^{\circ}$. We reproject the fisheye images in spherical projection, which have a vertical FoV of $189^{\circ}$ and a horizontal FoV of $120^{\circ}$. The lens projection is equidistant, and we use OpenCV with checkerboards to calibrate the camera parameters and the relative pose of the two cameras. Fig. 6 shows the qualitative results of our models compared to MODE. For both indoor and outdoor scenes, with models trained on 3D60 and Deep360 respectively, MCPDepth demonstrates noticeable improvements over MODE.
@@ 4.5 Ablation Study
Panorama Projection Tab. 4 demonstrates that cylindrical projection outperforms spherical projection in stereo matching, even with spherical convolutions (MODE [20]). Experimental results indicate that cylindrical projection is better suited for regular convolutions and more effective for panorama stereo matching.
@@
Circular Attention Tab. 5 shows that circular attention module effectively captures $360^{\circ}$ features and overcomes vertical axis distortion in panoramas. This module performs well across different projections and stereo-matching networks. For IGEV-Stereo [54], applying circular attention to the largest feature map (first scale) leads to significant performance gains.
@@ 5 Conclusion
We introduced MCPDepth, a two-stage framework for omnidirectional depth estimation via stereo matching between multiple cylindrical panoramas. We conducted both theoretical and experimental comparisons between spherical and cylindrical projections for stereo matching of panoramas and analyzed the advantages of cylindrical projection. The use of cylindrical projection maintains the linear epipolar constraint, similar to the Cassini projection, and also preserves the same definition of disparity in perspective images. Furthermore, it reduces distortion and provides the potential to apply stereo-matching models trained on perspective images to cylindrical panoramas. Additionally, cylindrical projection avoids using customized kernels, making it significantly simpler for deployment to embedded devices. Apart from this, the circular attention module effectively overcomes the distortion along the vertical axis of panoramas and captures $360^{\circ}$ features. The experimental results show that MCPDepth achieves state-of-the-art performance on the outdoor synthetic dataset Deep360 and the indoor real-scene dataset 3D60.
Supplementary Material
6 Framework
6.1 Stereo Matching
Fig. 8 shows the stereo matching network architecture with the circular attention module. Fig. 7 compares the attention heatmaps before and after circular attention. The attention module focuses primarily on the edges of objects and textured areas while paying less attention to the background and center areas of objects. This is because the disparities in the background are small, and the disparities in the center of objects are similar and contiguous.
Figure 7: Comparison of heat maps before and after circular attention module.
Figure 8: The architecture of the stereo matching network.
6.2 Depth Fusion
We use the depth fusion architecture from MODE [20], which takes depth maps, confidence maps, and reference panoramas. For Deep360, it uses 6 depth maps, 6 confidence maps, and 4 panoramas. For 3D60, it uses 3 of each. The Deep360 architecture is shown in Fig. 9.
Figure 9: Depth Fusion Architecture.
7 Training Details
During the stereo matching stage, we use 2 NVIDIA A40 GPUs to train our models with a batch size of 4 for the Deep360 dataset. Training takes 158 hours; For 3D60, we use a single NVIDIA A6000 GPU with batch size 4, which takes 252 hours to train. The model is trained for 45 epochs with a learning rate of 0.001, followed by a decay of the learning rate to 0.0001 for an additional 10 epochs. In the depth fusion stage, we train the network for 150 epochs with a learning rate of 0.0001.
8 Visualizations
Fig. 10 and Fig. 11 show the performance of stereo matching on Deep360 and 3D60 test datasets, our model can distinguish the foreground objects from the background, particularly in areas with severe distortion. Additionally, Fig. 12 illustrates the depth estimation performance on the 3D60 test dataset. With the help of more accurate disparity estimation, our method outperforms MODE on depth estimation performance even in areas where no ground truth is provided.
Figure 10: Disparity estimation results on the Deep360 test dataset compared to MODE.
Figure 11: Disparity estimation results on the 3D60 test dataset compared to MODE.
Figure 12: Qualitative comparisons of omnidirectional depth estimation methods on 3D60 compared with MODE.
9 Inference Time of Stereo Matching Models
The inference time of different stereo matching models on Deep360 and 3D60 datasets are shown in Tab. 6. The models are tested in both PyTorch and ONNX formats on NVIDIA GeForce RTX 3090. The panoramas have dimensions of $H\times W=1024\times 512$ on Deep360 and $H\times W=512\times 256$ on 3D60 for both cylindrical and spherical projections. For Deep360, the maximum disparity is set to 272 for cylindrical projection and 192 for spherical projection, while on 3D60, the maximum disparity is set to 256 for all projections. The inference time for the circular attention module is 18.22 ms for Deep360 and 2.81 ms for 3D60.
Our stereo matching model without circular attention outperforms MODE in stereo matching performance. Additionally, it exhibits a slightly shorter inference time when handling the same maximum disparity range, as demonstrated in Tab. 6 on the 3D60 dataset compared to MODE. Furthermore, our stereo matching model, while marginally slower than MODE, provides significant advantages when exporting to the model in ONNX format. Our model in ONNX format has a much shorter inference time than MODE in PyTorch format when handling the same maximum disparity range.
Table 6: Inference time of stereo matching model. ”CA” denotes Circular Attention.
Dataset Methods Projection PyTorch (ms) ONNX (ms)
Deep360 [20] MODE [20] Cassini 200.62 -
Ours w/o CA Cylindrical 238.61 197.63
Ours Cylindrical 266.27 226.30
3D60 [61] MODE [20] Cassini 66.74 -
Ours w/o CA Cylindrical 66.21 46.11
Ours Cylindrical 69.12 51.24

### v2

@@ 3.3 Circular Attention
Figure 4: (a) displays the circular attention module in the stereo matching network. (b) represents our attention applied along the vertical axis. $\bigoplus$ denotes element-wise sum. $\bigotimes$ denotes matrix multiplication. Blue boxes are $1\times 1$ convolution and orange boxes are relative positional encoding.
@@
To overcome vertical axis distortion and capture the circular $360{{}^{\circ}}$ features, we introduce a circular attention module. Conventional CNNs have limited receptive fields, which is restrictive for $360{{}^{\circ}}$ FoV panoramas. The circular attention module, placed between feature extraction and cost volume construction is flexible and can be easily integrated since it maintains the input dimension. Besides, it only calculates the relations along the vertical axis, conserving more computing costs compared to global self-attention approaches. Fig. 4 (a) demonstrates our circular attention module.
@@
where $\mathcal{N}_{1\times m}(o)$ is the local $1\times m$ region centered around location $o=(i,j)$. $r^{q}_{p-o}\in\mathbb{R}^{d_{q}}$ is the learnable relative positional encoding for queries and the inner product $q^{T}_{o}r^{q}_{p-o}$ measures the compatibility from location $p$ to location $o$. Similarly, the learnable vectors $r^{k}_{p-o}\in\mathbb{R}^{d_{q}}$ and $r^{v}_{p-o}\in\mathbb{R}^{d_{out}}$ are positional encodings for keys and values. Our circular attention reduces the computation to $(\mathcal{O}(hwm))$.
@@ 4.1 Datasets
We train and evaluate our framework on Deep360 [19] and 3D60 [65], which include outdoor and indoor scenes. We evaluate both stereo matching and depth estimation. For Deep360, four $360{{}^{\circ}}$ cameras are arranged horizontally in a square. Panoramas from all four views are used for evaluation. We use six stereo pairs for training and testing. For 3D60, three $360{{}^{\circ}}$ cameras are arranged vertically in an equilateral right triangle. Panoramas from two of three views are used for evaluation. The resolutions are $1024\times 512$ and $512\times 256$, respectively.
@@ 4.2 Evaluation Metrics
Following MODE [19], we evaluate stereo matching performance using MAE (mean absolute error), RMSE (root mean square error), Px1,3,5 (percentage of outliers with pixel error >> 1, 3, 5), D1 [30] (percentage of outliers with pixel error >> 3 and >> 5%). We evaluate depth estimation performance using MAE, RMSE, AbsRel (absolute relative error), SqRel (square relative error), SILog [7](scale-invariant logarithmic error), $\delta 1,2,3$ [16] (accuracy with threshold that $\max(\tfrac{\hat{y}}{y^{\star}},\tfrac{y^{\star}}{\hat{y}})<1.25,1.25^{2},1.25^{3}$).
@@ 4.3 Implementation Details
We apply nearest-neighbor interpolation for cylindrical/cubic disparity maps generalization and bilinear interpolation for cylindrical/cubic panoramas generalization, both derived from spherical inputs.
In the stereo matching stage, cylindrical panoramas have a $360{{}^{\circ}}$ vertical FoV and a horizontal FoV of less than $180{{}^{\circ}}$. We evaluate the central part of disparity maps in the Cassini domain with horizontal FoV = $2\arctan(\pi/2)\approx 105{{}^{\circ}}$ for both datasets. This FoV yields cylindrical panorama size equivalent to spherical. In the fusion stage, we evaluate the entire omnidirectional depth map with a $360{{}^{\circ}}$ horizontal and $180{{}^{\circ}}$ vertical FoV.
4.4 Experimental Results
Training on Perspective Images and Testing on Panoramas The pre-trained models of PSMNet [2] and IGEV-Stereo [56] are trained on Scene Flow [27], which contain only perspective images. CREStereo [17], trained on mixed datasets, exhibits better generalization. The performance of stereo matching with different projections on Deep360 is shown in Tab. 2. Acquiring panoramas and their depth ground truth is difficult, the experimental results demonstrate the potential to apply stereo-matching models trained on perspective images to cylindrical panoramas.
@@ 4.4 Experimental Results
For omnidirectional depth estimation, we compare our method with other multi-view omnidirectional depth estimation methods including UNiFuse [14], CSDNet [18], 360SD-Net [47], OmniMVS [52], and MODE [19]. We report the results from MODE. Tab. 4 shows that our method achieves an 18.8% MAE reduction on Deep360 and 19.9% on 3D60 compared to the previous best results, confirming its effectiveness for diverse real-world panoramas.
Table 3: Quantitative results of stereo matching methods on Deep360 and 3D60 test datasets. The top three results for each metric are highlighted with a first, second, and third background, respectively.
Dataset Method Projection Kernel Type MAE $\downarrow$ RMSE $\downarrow$ Px1 (%) $\downarrow$ Px3 (%) $\downarrow$ Px5 (%) $\downarrow$ D1 (%) $\downarrow$
Deep360 AANet [57] Cassini Regular 0.3427 1.5703 5.2050 2.1515 1.2847 1.9817
360SD-Net [47] Cassini Regular 0.5262 1.6459 3.8794 1.3389 0.8425 1.2989
PSMNet [2] Cassini Regular 0.2703 1.4790 3.3556 1.1979 0.7538 1.1708
MODE [19] Cassini Spherical 0.2309 1.4014 2.8801 1.0488 0.6562 1.0326
Ours Cylindrical Regular 0.2112 1.3903 2.5713 1.0009 0.6376 0.9828
3D60 MODE [19] Cassini Spherical 0.2258 0.5265 2.9441 0.6482 0.2978 0.6478
Ours Cylindrical Regular 0.1773 0.4654 2.2298 0.5282 0.2564 0.5279
@@
Dataset Method Kernel Type MAE $\downarrow$ RMSE $\downarrow$ AbsRel $\downarrow$ SqRel $\downarrow$ SILog $\downarrow$ $\delta 1\%\uparrow$ $\delta 2\%\uparrow$ $\delta 3\%\uparrow$
Deep360 OmniMVS [52] Regular 8.8865 59.3043 0.1073 2.9071 0.2434 94.9611 97.5495 98.2851
360SD-Net [47] Regular 11.2643 66.5789 0.0609 0.5973 0.2438 94.8594 97.2050 98.1038
CSDNet [18] Spherical 6.6548 36.5526 0.1553 1.7898 0.2475 86.0836 95.1589 97.7562
UniFuse [14] Regular 3.9193 28.8475 0.0546 0.3125 0.1508 96.0269 98.2679 98.9909
MODE [19] Spherical 3.2483 24.9391 0.0365 0.0789 0.1104 97.9636 99.0987 99.4683
Ours+Cubic Regular 5.0309 36.1907 0.0785 0.4410 0.1781 94.5960 98.1782 98.9406
Ours Regular 2.6384 21.6692 0.0304 0.1153 0.1033 98.2557 99.2101 99.5227
3D60 360SD-Net [47] Regular 0.0762 0.2639 0.0300 0.0117 1.4578 97.6751 98.6603 99.0417
CSDNet [18] Spherical 0.2067 0.4225 0.0908 0.0241 0.1273 91.9537 98.3936 99.5109
UniFuse [14] Regular 0.1868 0.3947 0.0799 0.0246 0.1126 93.2860 98.4839 99.4828
MODE [19] Spherical 0.0713 0.2631 0.0224 0.0031 0.0512 99.1283 99.7847 99.9250
Ours Regular 0.0571 0.1903 0.0199 0.0027 0.0401 99.3933 99.8506 99.9418
Fig. 5 illustrates the superior performance on Deep360, effectively handling severe distortions while preserving finer object details and the edges between the foreground and background better than MODE.
@@
Performance on Real Scenarios We evaluate our models on real-world fisheye image pairs. We reproject the fisheye images in Cassini projection, which have a vertical FoV of $189{{}^{\circ}}$ and a horizontal FoV of $120^{\circ}$. The lens projection is equidistant, and we use OpenCV with checkerboards to calibrate the camera parameters and the relative pose of the two cameras.  Fig. 6 shows the qualitative results of our models compared to MODE. For both indoor and outdoor scenes, with models trained on 3D60 and Deep360 respectively, MCPDepth demonstrates noticeable improvements over MODE, particularly in the highly distorted areas.
@@ 4.5 Ablation Study
Panorama Projection Tab. 5 demonstrates that cylindrical projection significantly outperforms spherical projection in stereo matching, even when applying spherical convolutions on spherical panoramas (MODE). Furthermore, we compare the performance of different projections on depth estimation in Tab. 4, demonstrating that cylindrical projection is the most suitable projection for regular convolutions, making it more effective for panorama stereo matching and depth estimation. These benefits may extend to other panorama-related vision tasks.
Circular Attention Tab. 6 shows that, although designed to mitigate vertical distortion for cylindrical projection, our circular attention module consistently improves performance across various panoramic projections and stereo-matching networks. This lightweight module delivers significant accuracy gains with minimal additional computation, evaluated in the Cassini domain for spherical/cylindrical panoramas and the cubic domain for cubic panoramas. For IGEV-Stereo [56], applying circular attention to the largest feature map (first scale) yields substantial performance improvements.
@@ 5 Conclusion
We present MCPDepth, a novel two-stage framework for omnidirectional depth estimation through stereo matching from multiple cylindrical panoramas. Our comprehensive theoretical and experimental comparisons on different panoramic projections highlight the distinct advantages of cylindrical projection. Cylindrical projection maintains the linear epipolar constraint and preserves the definition of disparity as in perspective images. It effectively reduces distortion, enabling the application of stereo-matching models trained on perspective images to cylindrical panoramas. Additionally, cylindrical projection eliminates the need for customized kernels, simplifying deployment on embedded devices. Our circular attention module addresses vertical-axis distortions in cylindrical panoramas and captures $360{{}^{\circ}}$ features, and can be extended to other projections. Experimental results demonstrate that MCPDepth achieves state-of-the-art performance on both the outdoor dataset Deep360 and the indoor dataset 3D60.
Limitations Cylindrical panoramas are limited in their ability to capture a full $180{{}^{\circ}}$ horizontal FoV. As a result, at least 3 cameras are required to ensure complete coverage. Future work should explore optimizing the horizontal FoV of cylindrical panoramas to balance performance and computational resources.
Acknowledgments
We gratefully acknowledge the advanced computational resources provided by Engineering IT and Research Infrastructure Services at Washington University in St. Louis.